# The Impact of Quantization on Retrieval-Augmented Generation: An Analysis of Small LLMs

### v2

@@ Abstract.
Retrieval Augmented Generation, Quantization, Efficiency, Large Language Models, Personalization
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Personalizationccs: Computing methodologies Natural language generation
1. Introduction
Large Language Model (LLM) outputs can be enhanced by fetching relevant documents via a retriever and adding them as context for the prompt. The LLM can generate an output grounded with relevant information with the added context. This process is called Retrieval Augmented Generation (RAG). RAG has many benefits such as improving effectiveness in downstream tasks (Huang et al., 2023; Ma et al., 2023; Shi et al., 2023; Xu et al., 2023), reducing hallucinations (Proser, [n. d.]), increasing factuality (Nakano et al., 2022), by-passing knowledge cut-offs, and presenting proprietary data that is not available to the LLMs.
The performance of RAG depends on the number, quality, and relevance of the retrieved documents (Gao et al., 2024). To perform RAG, many tasks demand a lot of passages extracted from multiple, unstructured documents: For question-answering tasks, the answer might be scattered around many documents because of ambiguity or the time-series nature of the question (eg. price change of a stock). For more open-ended tasks like personalization, many documents from different sources might be needed to capture the characteristics of the individual. Therefore to handle RAG in these tasks, an LLM needs to look at multiple sources, identify the relevant parts, and compose the most plausible answer (Gao et al., 2024).
LLMs do not pay the same attention to their whole context windows, meaning the placement of documents in the prompt directly affects the final output (Liu et al., 2023). On top of that, some of the retrieved documents may be unrelated to the task, or they may contain contradictory information compared to the parametric knowledge of the LLM (Xu et al., 2024). An LLM has to overcome these challenges to leverage RAG to its advantage. Xu et al. (2023) have shown that an open-source 70B LLM (Touvron et al., 2023) equipped with RAG can beat proprietary models, meaning it is not necessary to use an LLM in the caliber of GPT-4 (OpenAI, 2023) to implement RAG. Still, for many use cases, it might not be feasible to deploy a 70B LLM as it is computationally demanding. To decrease the computational demand of LLMs, post-training quantization can be used. Quantization drastically reduces the required amount of RAM to load a model and can increase the inference speed by more than 3 times (Dettmers et al., 2023; Frantar et al., 2023). Despite the benefits, quantization affects LLMs differently depending on their size (Li et al., 2024). For capabilities that are important to RAG, such as long-context reasoning, smaller LLMs (¡13B) are found to be more sensitive to quantization (Li et al., 2024).
Figure 1. Prompts used for both datasets. The ones on the top represent $k=0$ (zero-shot, no retrieved documents) and the ones on the bottom are for $k>0$ settings (RAG). The green text is the model output. Line endings are not shown for space reasons.
@@ 1. Introduction
Prompts used during experiments. LaMP-3U prompts are on the left and LaMP-5U prompts are on the right
@@
In this paper, we investigate the effectiveness of quantization on RAG-enhanced 7B and 8B LLMs. We evaluate the full (FP16) and quantized (INT4) versions of multiple LLMs on two personalization tasks taken from the LaMP (Salemi et al., 2023) benchmark. To better study how quantized LLMs perform in longer contexts, we compared the performance gap between FP16 and INT4 models with an increasing number of retrieved documents. We chose personalization because it is a challenging task to perform with RAG as it demands long-context reasoning over many documents. Contrary to question-answering where the LLM has to find the correct answer from a couple of documents, personalization requires the LLM to carefully study a person’s style from all the provided documents. Our findings show that the effect of quantization depends on the model and the task: we find almost no drop in performance for OpenChat while LLaMA2 seems to be more sensitive. Our experiments show that quantized smaller LLMs can be good candidates for RAG pipelines, especially if efficiency is essential.
2. Approach
2.1. LLMs
@@
Starting with LLaMA2-7B (Chat-hf) (Touvron et al., 2023) to have a baseline, we experiment with the following LLMs: LLaMA3-8B (Meta, 2024), Zephyr (Beta) (Tunstall et al., 2023), OpenChat (3.5) (Wang et al., 2023), and Starling (LM-alpha) (Zhu et al., 2023). These models were chosen because they were the highest-ranked 7B and 8B LLMs in the Chatbot Arena Leaderboard (Zheng et al., 2023) according to the Elo ratings at the time of writing. Since all models except LLaMA are finetuned variants of Mistral-7B (Jiang et al., 2023), we add Mistral-7B (Instruct-0.1) [^1] Although there is an updated v0.2 version of Mistral-7B, we used v0.1 to match the other LLMs that are finetuned on it to our experiments too. We use Activation-aware Weight Quantization (AWQ) as it outperforms other methods (Lin et al., 2023).
2.2. Tasks and Datasets
@@
We use the LaMP benchmark that offers 7 personalization datasets with either a classification or a generation task (Salemi et al., 2023). To represent both types of tasks, we chose one dataset from each: LaMP-3 (“Personalized Product Rating”) and LaMP-5 (“Personalized Scholarly Title Generation”). LaMP-3 is composed of product reviews and their corresponding scores. For each user, one of the review–score pairs is chosen as the target and other pairs become the user profile. The LLM’s task, in this case, is to predict the score given a review using the other review–score pairs of the same user. LaMP-5 aims to generate a title for an academic paper based on the abstract. In this case, the user profile consists of abstract–title pairs that demonstrate the writing style of the user (scholar). The task of the LLM is to generate a title for the given abstract by incorporating the writing style of the scholar. Those datasets were chosen because compared to the other ones, on average, they had more samples in their user profiles, and the samples were longer. Therefore, they represented a better opportunity to evaluate RAG effectiveness as the retrieval part would be trickier.
@@
We work with the user-based splits (LaMP-3U, LaMP-5U) where the user appears only in one of the data splits (Salemi et al., 2023). The labels for the test sets are not publicly available (results can be obtained by submitting the predictions to the leaderboard) and since we did not fine-tune our models, we chose to use the validation sets for evaluation. For both datasets, we noticed that some samples do not fit in the context windows. After analyzing the overall length of the samples, we concluded that those cases only represent a tiny minority and removed data points that are not in the 0.995th percentile. For LaMP-5U, we also removed abstracts that consisted only of the text “no abstract available”. There are 2500 samples in the validation sets, and we have 2487 samples left after the preprocessing steps for both datasets.
@@
Evaluation. We used mean absolute error (MAE) for LaMP-3 and Rouge-L (Lin, 2004) for LaMP-5, following the LaMP paper (Salemi et al., 2023). Their experiments also include root mean square error (RMSE) and Rouge-1 scores, but we found that the correlation between MAE and RMSE is 0.94, and between Rouge-1 and Rouge-L is 0.99. Therefore, we do not include those metrics in our results. The prompts we use are shown in Figure 1. Even though the LLMs are instructed to output only the score or the title, we notice that some are prone to give lengthy answers such as “Sure, here is the title for the given abstract, Title: (generated title)”. We apply a post-processing step on the LLM outputs to extract only the score or the title before evaluation.
2.3. Retrieval
@@
We conduct the experiments with the following number of retrieved documents: $k\in\{0,1,3,5,max\_4K,max\_8K\}$. 00 refers to zero-shot without any retrieval and $max$ is the maximum number of documents that can be put into the prompt, given the context window of the LLM. LLaMA2 has a context window of 4096 tokens while other models have 8192 tokens. To make it fair, we include two options for the $max$ setting: 4K and 8K. For $max\_4K$, we assume that all models have a 4096 token context window, we use the original 8192 token context windows for $max\_8K$. Consequently, LLaMA2-7B is not included in the $max\_8K$ experiments. To put it into perspective, the number of retrieved documents varies between 15-18 for $max\_4K$, and between 25-28 for $max\_8K$ in LaMP-5U, depending on the average length of documents in the user profile. As retrievers, we evaluate BM25 (Robertson et al., 1994) (BM25 Okapi)  [^2] https://pypi.org/project/rank-bm25/, Contriever (Izacard et al., 2022) (finetuned on MS-Marco), and DPR (Karpukhin et al., 2020) (finetuned on Natural Questions) [^3] https://huggingface.co/facebook/dpr-question_encoder-single-nq-base. Since we focus on efficiency by reducing the computational load, the retrievers are not finetuned on the datasets.
@@
Table 1. The absolute percentage change between FP16 and INT4 scores, using Contriever. More than a 5% drop in performance is highlighted in red. For MAE, the lower is better while the inverse is true for Rouge-L.
[…]
Table showing the results of the experiments for Contriever. We list the difference in scores between the FP16 and INT4 variants of the models by how much the performance changes in percentages.
3. Results
@@
Figure 2. Results for both datasets. The upper and lower borders of each colored area represent the quantized and not-quantized performances of the models, and the corresponding lines are the mean of both.
[…]
Results for both datasets using 3 different retrievers. The left plot shows the MAE for the third dataset and the right plot shows the Rouge-L score. The x-axis is composed of the values of the number of retrieved documents. It shows the same results as Table 1.
LLMs. The results are shown in Table 1. We see the dominance of OpenChat in both datasets. Zephyr performs very close to OpenChat in LaMP-3U but falls far behind in LaMP-5U. The same can be said for Starling but reversed. Mistral-7B performs stable in both datasets albeit being slightly behind OpenChat. Overall, LLaMA2 performs the worst as it is below average in both datasets. Despite being the dominant small LLM currently, LLaMA3 is not the best for both tasks, despite performing reasonably well in LaMP-3U. Interestingly, LLaMA3 has the best zero-shot score in LaMP-5U but it struggles to improve itself with retrieval.
Quantization. How much an LLM gets affected by quantization seems to be related to how well it performs the task. OpenChat suffers almost no performance degradation from quantization. On the contrary, LLaMA2 seems very sensitive, especially when the number of retrieved documents is increased. Starling suffers no significant consequence from quantization in LaMP-5U where it performs well, but it does suffer in LaMP-3U. There also seems to be a disparity between the tasks as quantized LLMs perform much worse in LaMP-3U than in LaMP-5U.
Number of retrieved documents. Figure 2b shows that LLM performance is saturated with a couple of documents, and the improvement obtained from more is marginal. In LaMP-5U, adding more than 5 documents starts to hurt the performance: Figure 2b shows an inverse-U-shaped distribution for all LLMs except LLaMA3. For some models, performance even drops below the zero-shot setting when all the available context window is filled with retrieved documents. The LaMP-3U results (Figure 2a) continue to improve after adding more than 5 documents as can be observed from $max\_4K$ scores, but MAE also starts to get worse for $max\_8K$.
@@
Retrievers. Figure 2b shows that the three retrievers gave almost identical results, albeit BM25 being marginally behind the others. Also in LaMP-5U, the gap between the FP16 and INT4 LLaMA2 varies slightly. Other than that, the retriever model does not have a noticeable impact on the personalization tasks we experimented with. The patterns we found regarding LLMs, quantization, and the number of retrieved documents are the same for all the retrievers.
@@
ChatGPT (Salemi et al., 2023) 0.658 0.336 ?
FlanT5-XXL (Salemi et al., 2023) 0.282 0.424 43 GB
@@
Table 2. Our results compared with LaMP. Results indicated with * are not significantly lower than the reported best result (FlanT5-XXL). A quantized 7B LLM can perform on par with a larger model while being much more efficient.
[…]
Table showing the Rouge-L scores obtained from the experiments
Benchmark comparison. Finally, we compared our findings with the RAG results from LaMP (Salemi et al., 2023). Table 2 shows that OpenChat (Contriever, $k=5$) can beat FlanT5-XXL in LaMP-3U and performs very close in LaMP-5U. More importantly, its quantized counterpart has very similar results. Since we do not have the per-sample scores for the baseline models from LaMP, we perform a one-sample t-test on the Rouge-L scores. The corresponding $p$ value of 0.29 shows a non-significant difference between the results. Moreover, the results from the LaMP paper are with finetuned retrievers while our results are with non-finetuned retrievers. This indicates that a quantized-7B LLM can compete and even outperform a bigger model on personalization with RAG.
@@
4. Discussion
Our results show that some LLMs (in particular OpenChat) can be successful in RAG pipelines, even after quantization, but the performance is LLM- and task-dependent. The method of quantization affects LLMs differently (Li et al., 2024). Thus, the relationship between quantization and RAG performance is not straightforward and can be studied more extensively. Still, our results indicate that when a small LLM performs the task well, its AWQ-quantized counterpart performs on par.
The differing performance of some LLMs between the datasets may be partly due to prompting. LLMs are sensitive to prompts, and a prompt that works for one LLM may not work for another one (Sclar et al., 2023). The most peculiar result is the lackluster performance of LLaMA3 in LaMP-5U. LLaMA3 is a recently released model trained with an extensive pretraining corpus (Meta, 2024). It has a higher chance of seeing the abstracts presented in the LaMP-5U in its pretraining data. This may explain its superior zero-shot performance. LLMs suffer from a knowledge conflict between their parametric information and the contextual information presented through retrieval (Xu et al., 2024). If LLaMA3 had already memorized some of the titles of the abstracts in LaMP-5U, it might result in a knowledge conflict when similar abstract-title pairs of the same author are presented. This may explain the reduced improvement in its performance with retrieval.
LLMs have been shown to struggle with too many retrieved documents (Liu et al., 2023), and our findings are in accordance. Our results indicate that more than $>5$ documents do not help and can even hurt performance. From prior studies, we know that LLMs focus more on the bottom and the top of their context window (Liu et al., 2023). We progressively put the most relevant documents starting from the bottom to the top. Therefore especially for $k=max$ settings, the less relevant documents are put on the top. This situation might hurt the LLM performance as it would focus on the most and the least related information in this case. That being said, state-of-the-art LLMs with more than 7B parameters also suffer from the same phenomenon even when not quantized (Liu et al., 2023). Although quantization increases the risk of worsened long-context performance, we cannot conclude that it is the sole perpetrator, as this is an inherent problem for all LLMs.
5. Conclusion

### v3

@@ Abstract
keywords:
@@
1 Introduction
Large Language Model (LLM) outputs can be enhanced by fetching relevant documents via a retriever and adding them as context for the prompt. The LLM can generate an output grounded with relevant information with the added context. This process is called Retrieval Augmented Generation (RAG). RAG has many benefits such as improving effectiveness in downstream tasks [1, 2, 3, 4], reducing hallucinations [5], increasing factuality [6], by-passing knowledge cut-offs, and presenting proprietary data that is not available to the LLMs.
The performance of RAG depends on the number, quality, and relevance of the retrieved documents [7]. To perform RAG, many tasks demand a lot of passages extracted from multiple, unstructured documents: For question-answering tasks, the answer might be scattered around many documents because of ambiguity or the time-series nature of the question (eg. price change of a stock). For more open-ended tasks like personalization, many documents from different sources might be needed to capture the characteristics of the individual. Therefore to handle RAG in these tasks, an LLM needs to look at multiple sources, identify the relevant parts, and compose the most plausible answer [7].
LLMs do not pay the same attention to their whole context windows, meaning the placement of documents in the prompt directly affects the final output [8]. On top of that, some of the retrieved documents may be unrelated to the task, or they may contain contradictory information compared to the parametric knowledge of the LLM [9]. An LLM has to overcome these challenges to leverage RAG to its advantage. Xu et al. [4] have shown that an open-source 70B LLM [10] equipped with RAG can beat proprietary models, meaning it is not necessary to use an LLM in the caliber of GPT-4 [11] to implement RAG. Still, for many use cases, it might not be feasible to deploy a 70B LLM as it is computationally demanding. To decrease the computational demand of LLMs, post-training quantization can be used. Quantization drastically reduces the required amount of RAM to load a model and can increase the inference speed by more than 3 times [12, 13]. Despite the benefits, quantization affects LLMs differently depending on their size [14]. For capabilities that are important to RAG, such as long-context reasoning, smaller LLMs (<13B) are found to be more sensitive to quantization [14].
Figure 1: Prompts used for both datasets. The ones on the top represent $k=0$ (zero-shot, no retrieved documents) and the ones on the bottom are for $k>0$ settings (RAG). The green text is the model output. Line endings are not shown for space reasons.
@@ 1 Introduction
In this paper, we investigate the effectiveness of quantization on RAG-enhanced 7B and 8B LLMs. We evaluate the full (FP16) and quantized (INT4) versions of multiple LLMs on two personalization tasks taken from the LaMP [15] benchmark. To better study how quantized LLMs perform in longer contexts, we compared the performance gap between FP16 and INT4 models with an increasing number of retrieved documents. We chose personalization because it is a challenging task to perform with RAG as it demands long-context reasoning over many documents. Contrary to question-answering where the LLM has to find the correct answer from a couple of documents, personalization requires the LLM to carefully study a person’s style from all the provided documents. Our findings show that the effect of quantization depends on the model and the task: we find almost no drop in performance for OpenChat while LLaMA2 seems to be more sensitive. Our experiments show that quantized smaller LLMs can be good candidates for RAG pipelines, especially if efficiency is essential.
2 Approach
2.1 LLMs
@@
Starting with LLaMA2-7B (Chat-hf) [10] to have a baseline, we experiment with the following LLMs: LLaMA3-8B [16], Zephyr (Beta) [17], OpenChat (3.5) [18], and Starling (LM-alpha) [19]. These models were chosen because they were the highest-ranked 7B and 8B LLMs in the Chatbot Arena Leaderboard [20] according to the Elo ratings at the time of writing. Since all models except LLaMA are finetuned variants of Mistral-7B [21], we add Mistral-7B (Instruct-0.1) [^1] Although there is an updated v0.2 version of Mistral-7B, we used v0.1 to match the other LLMs that are finetuned on it to our experiments too. We use Activation-aware Weight Quantization (AWQ) as it outperforms other methods [22].
2.2 Tasks and Datasets
@@
We use the LaMP benchmark that offers 7 personalization datasets with either a classification or a generation task [15]. To represent both types of tasks, we chose one dataset from each: LaMP-3 (“Personalized Product Rating”) and LaMP-5 (“Personalized Scholarly Title Generation”). LaMP-3 is composed of product reviews and their corresponding scores. For each user, one of the review–score pairs is chosen as the target and other pairs become the user profile. The LLM’s task, in this case, is to predict the score given a review using the other review–score pairs of the same user. LaMP-5 aims to generate a title for an academic paper based on the abstract. In this case, the user profile consists of abstract–title pairs that demonstrate the writing style of the user (scholar). The task of the LLM is to generate a title for the given abstract by incorporating the writing style of the scholar. Those datasets were chosen because compared to the other ones, on average, they had more samples in their user profiles, and the samples were longer. Therefore, they represented a better opportunity to evaluate RAG effectiveness as the retrieval part would be trickier.
@@
We work with the user-based splits (LaMP-3U, LaMP-5U) where the user appears only in one of the data splits [15]. The labels for the test sets are not publicly available (results can be obtained by submitting the predictions to the leaderboard) and since we did not fine-tune our models, we chose to use the validation sets for evaluation. For both datasets, we noticed that some samples do not fit in the context windows. After analyzing the overall length of the samples, we concluded that those cases only represent a tiny minority and removed data points that are not in the 0.995th percentile. For LaMP-5U, we also removed abstracts that consisted only of the text “no abstract available”. There are 2500 samples in the validation sets, and we have 2487 samples left after the preprocessing steps for both datasets.
2.2.1 Evaluation
@@
We used mean absolute error (MAE) for LaMP-3 and Rouge-L [23] for LaMP-5, following the LaMP paper [15]. Their experiments also include root mean square error (RMSE) and Rouge-1 scores, but we found that the correlation between MAE and RMSE is 0.94, and between Rouge-1 and Rouge-L is 0.99. Therefore, we do not include those metrics in our results. The prompts we use are shown in Figure 1. Even though the LLMs are instructed to output only the score or the title, we notice that some are prone to give lengthy answers such as “Sure, here is the title for the given abstract, Title: (generated title)”. We apply a post-processing step on the LLM outputs to extract only the score or the title before evaluation.
2.3 Retrieval
@@
We conduct the experiments with the following number of retrieved documents: $k\in\{0,1,3,5,max\_4K,max\_8K\}$. 00 refers to zero-shot without any retrieval and $max$ is the maximum number of documents that can be put into the prompt, given the context window of the LLM. LLaMA2 has a context window of 4096 tokens while other models have 8192 tokens. To make it fair, we include two options for the $max$ setting: 4K and 8K. For $max\_4K$, we assume that all models have a 4096 token context window, we use the original 8192 token context windows for $max\_8K$. Consequently, LLaMA2-7B is not included in the $max\_8K$ experiments. To put it into perspective, the number of retrieved documents varies between 15-18 for $max\_4K$, and between 25-28 for $max\_8K$ in LaMP-5U, depending on the average length of documents in the user profile. As retrievers, we evaluate BM25 [24] (BM25 Okapi)  [^2] https://pypi.org/project/rank-bm25/, Contriever [25] (finetuned on MS-Marco), and DPR [26] (finetuned on Natural Questions) [^3] https://huggingface.co/facebook/dpr-question_encoder-single-nq-base. Since we focus on efficiency by reducing the computational load, the retrievers are not finetuned on the datasets.
Table 1: The absolute percentage change between FP16 and INT4 scores, using Contriever. More than a 5% drop in performance is highlighted in red. For MAE, the lower is better while the inverse is true for Rouge-L.
@@
3 Results
@@
Figure 2: Results for both datasets. The upper and lower borders of each colored area represent the quantized and not-quantized performances of the models, and the corresponding lines are the mean of both.
3.1 LLMs
The results are shown in Table 1. We see the dominance of OpenChat in both datasets. Zephyr performs very close to OpenChat in LaMP-3U but falls far behind in LaMP-5U. The same can be said for Starling but reversed. Mistral-7B performs stable in both datasets albeit being slightly behind OpenChat. Overall, LLaMA2 performs the worst as it is below average in both datasets. Despite being the dominant small LLM currently, LLaMA3 is not the best for both tasks, despite performing reasonably well in LaMP-3U. Interestingly, LLaMA3 has the best zero-shot score in LaMP-5U but it struggles to improve itself with retrieval.
3.2 Quantization
How much an LLM gets affected by quantization seems to be related to how well it performs the task. OpenChat suffers almost no performance degradation from quantization. On the contrary, LLaMA2 seems very sensitive, especially when the number of retrieved documents is increased. Starling suffers no significant consequence from quantization in LaMP-5U where it performs well, but it does suffer in LaMP-3U. There also seems to be a disparity between the tasks as quantized LLMs perform much worse in LaMP-3U than in LaMP-5U.
3.3 Number of retrieved documents
Figure 2 shows that LLM performance is saturated with a couple of documents, and the improvement obtained from more is marginal. In LaMP-5U, adding more than 5 documents starts to hurt the performance: Figure 2b shows an inverse-U-shaped distribution for all LLMs except LLaMA3. For some models, performance even drops below the zero-shot setting when all the available context window is filled with retrieved documents. The LaMP-3U results (Figure 2a) continue to improve after adding more than 5 documents as can be observed from $max\_4K$ scores, but MAE also starts to get worse for $max\_8K$.
@@ 3.3 Number of retrieved documents
3.4 Retrievers
Figure 2 shows that the three retrievers gave almost identical results, albeit BM25 being marginally behind the others. Also in LaMP-5U, the gap between the FP16 and INT4 LLaMA2 varies slightly. Other than that, the retriever model does not have a noticeable impact on the personalization tasks we experimented with. The patterns we found regarding LLMs, quantization, and the number of retrieved documents are the same for all the retrievers.
Table 2: Our results compared with LaMP. Results indicated with * are not significantly lower than the reported best result (FlanT5-XXL). A quantized 7B LLM can perform on par with a larger model while being much more efficient.
@@ 3.4 Retrievers
ChatGPT [15] 0.658 0.336 ?
FlanT5-XXL [15] 0.282 0.424 43 GB
@@
3.5 Benchmark comparison
Finally, we compared our findings with the RAG results from LaMP [15]. Table 2 shows that OpenChat (Contriever, $k=5$) can beat FlanT5-XXL in LaMP-3U and performs very close in LaMP-5U. More importantly, its quantized counterpart has very similar results. Since we do not have the per-sample scores for the baseline models from LaMP, we perform a one-sample t-test on the Rouge-L scores. The corresponding $p$ value of 0.29 shows a non-significant difference between the results. Moreover, the results from the LaMP paper are with finetuned retrievers while our results are with non-finetuned retrievers. This indicates that a quantized-7B LLM can compete and even outperform a bigger model on personalization with RAG.
@@ 3.5 Benchmark comparison
4 Discussion
Our results show that some LLMs (in particular OpenChat) can be successful in RAG pipelines, even after quantization, but the performance is LLM- and task-dependent. The method of quantization affects LLMs differently [14]. Thus, the relationship between quantization and RAG performance is not straightforward and can be studied more extensively. Still, our results indicate that when a small LLM performs the task well, its AWQ-quantized counterpart performs on par.
The differing performance of some LLMs between the datasets may be partly due to prompting. LLMs are sensitive to prompts, and a prompt that works for one LLM may not work for another one [27]. The most peculiar result is the lackluster performance of LLaMA3 in LaMP-5U. LLaMA3 is a recently released model trained with an extensive pretraining corpus [16]. It has a higher chance of seeing the abstracts presented in the LaMP-5U in its pretraining data. This may explain its superior zero-shot performance. LLMs suffer from a knowledge conflict between their parametric information and the contextual information presented through retrieval [9]. If LLaMA3 had already memorized some of the titles of the abstracts in LaMP-5U, it might result in a knowledge conflict when similar abstract-title pairs of the same author are presented. This may explain the reduced improvement in its performance with retrieval.
LLMs have been shown to struggle with too many retrieved documents [8], and our findings are in accordance. Our results indicate that more than $>5$ documents do not help and can even hurt performance. From prior studies, we know that LLMs focus more on the bottom and the top of their context window [8]. We progressively put the most relevant documents starting from the bottom to the top. Therefore especially for $k=max$ settings, the less relevant documents are put on the top. This situation might hurt the LLM performance as it would focus on the most and the least related information in this case. That being said, state-of-the-art LLMs with more than 7B parameters also suffer from the same phenomenon even when not quantized [8]. Although quantization increases the risk of worsened long-context performance, we cannot conclude that it is the sole perpetrator, as this is an inherent problem for all LLMs.
5 Conclusion